# Logarithmic delocalization of low temperature 3D Ising and Potts interfaces above a hard floor

### v1

@@ 4.2. Exact additivity of the rates
We define certain high probability cone events in the random cluster model, which we will show all have probability $1-\varepsilon_{\beta}$ conditional on $A_{\mathsf{o},h_{1}+h_{2}},A_{\mathsf{o},h_{1}},A_{\mathsf{o},h_{2}}$ respectively in $\widetilde{\mathbf{P}}^{\mathsf{red}}_{n}$. We then show the approximate additivity of the rates under those events. The events are defined as follows:
@@
Note that conditioning on $\mathcal{E}_{\sqcup}$ is equivalent to conditioning on the aforementioned five plaquettes to be dual-to-closed, as well as on the plaquettes that are 1-connected to them and have height $<0$ to be dual-to-open. Let $\zeta$ be the FK–Potts boundary condition with these dual-to-open and dual-to-closed plaquettes, the wired $\mathsf{red}$ boundary at $\partial\Lambda_{n,n}^{\prime}$, and $\sigma_{\mathsf{o}-{\mathfrak{e}_{3}}}=\mathsf{blue}$ (by symmetry, any $\mathsf{nonR}$ color is equivalent). Let $F$ be the set of plaquettes fixed by $\zeta$. Then we can write
@@
Now, we observe that the random cluster marginal of ${\mathbf{P}}^{\eta^{\prime}}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}$ has the same law as the random cluster measure on $\Lambda_{n,n}^{\prime}\setminus E^{\prime}$ with all the dual-to-open and dual-to-closed plaquettes in $\eta^{\prime}$ as a boundary condition, further conditional on the event that there is no $\omega$-path from $\mathsf{o}-{\mathfrak{e}_{3}}$ to $\partial\Lambda_{n,n}^{\prime}$. Call this random cluster measure $\pi^{\eta^{\prime}(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, where ${\mathfrak{D}}_{\mathsf{o}}$ denotes this separation event for $\mathsf{o}-{\mathfrak{e}_{3}}$. Furthermore, by the FK–Potts coupling, the joint configuration under ${\mathbf{P}}^{\eta^{\prime}}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}$ can be sampled by first sampling a random cluster configuration $\omega$ under $\pi^{\eta^{\prime}(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, assigning vertices in the same $\omega$-component of $\mathsf{o}-{\mathfrak{e}_{3}}$ the same $\mathsf{nonR}$ color as given by $\eta^{\prime}$ (by symmetry, we may assume this is $\mathsf{blue}$), assigning vertices in the $\omega$-component of $\partial\Lambda_{n,n}^{\prime}$ the color $\mathsf{red}$, and assigning all other $\omega$-components colors chosen independently at random. In the same way (with analogous notation), we can sample a joint configuration under ${\mathbf{P}}^{\zeta}_{\Lambda_{n,n}^{\prime}\setminus F}$ by first sampling a random cluster configuration $\omega$ under $\pi^{\zeta(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus F}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, assigning vertices in the same $\omega$-component of $\mathsf{o}-{\mathfrak{e}_{3}}$ the color $\mathsf{blue}$, assigning vertices in the $\omega$-component of $\partial\Lambda_{n,n}^{\prime}$ the color $\mathsf{red}$, and assigning all other $\omega$-components colors chosen independently at random.
@@ Appendix A Generalized cluster expansion
The integral above is a constant, so it and the term $c^{r}(\tfrac{q}{2})^{r/c}$ can be abosrbed into the term $e^{-\beta r/c_{2}}$ by increasing the constant $c_{2}$ to some sufficiently large (but $\beta$ independent) constant $K$. ∎
@@ Acknowledgments
The research of R.G. is supported in part by NSF DMS-2246780. The research of E.L. was supported by the NSF grant DMS-2054833.

### v2

@@ 4.2. Exact additivity of the rates
We define certain high probability cone events in the random-cluster model, which we will show all have probability $1-\varepsilon_{\beta}$ conditional on $A_{\mathsf{o},h_{1}+h_{2}},A_{\mathsf{o},h_{1}},A_{\mathsf{o},h_{2}}$ respectively in $\widetilde{\mathbf{P}}^{\mathsf{red}}_{n}$. We then show the approximate additivity of the rates under those events. The events are defined as follows:
@@
Note that conditioning on $\mathcal{E}_{\sqcup}$ is equivalent to conditioning on the aforementioned five plaquettes to be dual-to-closed, as well as on the plaquettes that are 1-connected to them and have height <0<0 to be dual-to-open. Let $\zeta$ be the FK–Potts boundary condition with these dual-to-open and dual-to-closed plaquettes, the wired $\mathsf{red}$ boundary at $\partial\Lambda_{n,n}^{\prime}$, and $\sigma_{\mathsf{o}-{\mathfrak{e}_{3}}}=\mathsf{blue}$ (by symmetry, any $\mathsf{nonR}$ color is equivalent). Let $F$ be the set of plaquettes fixed by $\zeta$. Then we can write
@@
Now, we observe that the random-cluster marginal of ${\mathbf{P}}^{\eta^{\prime}}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}$ has the same law as the random-cluster measure on $\Lambda_{n,n}^{\prime}\setminus E^{\prime}$ with all the dual-to-open and dual-to-closed plaquettes in $\eta^{\prime}$ as a boundary condition, further conditional on the event that there is no $\omega$-path from $\mathsf{o}-{\mathfrak{e}_{3}}$ to $\partial\Lambda_{n,n}^{\prime}$. Call this random-cluster measure $\pi^{\eta^{\prime}(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, where ${\mathfrak{D}}_{\mathsf{o}}$ denotes this separation event for $\mathsf{o}-{\mathfrak{e}_{3}}$. Furthermore, by the FK–Potts coupling, the joint configuration under ${\mathbf{P}}^{\eta^{\prime}}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}$ can be sampled by first sampling a random-cluster configuration $\omega$ under $\pi^{\eta^{\prime}(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus E^{\prime}}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, assigning vertices in the same $\omega$-component of $\mathsf{o}-{\mathfrak{e}_{3}}$ the same $\mathsf{nonR}$ color as given by $\eta^{\prime}$ (by symmetry, we may assume this is $\mathsf{blue}$), assigning vertices in the $\omega$-component of $\partial\Lambda_{n,n}^{\prime}$ the color $\mathsf{red}$, and assigning all other $\omega$-components colors chosen independently at random. In the same way (with analogous notation), we can sample a joint configuration under ${\mathbf{P}}^{\zeta}_{\Lambda_{n,n}^{\prime}\setminus F}$ by first sampling a random-cluster configuration $\omega$ under $\pi^{\zeta(\mathsf{RC})}_{\Lambda_{n,n}^{\prime}\setminus F}(\cdot\mid{\mathfrak{D}}_{\mathsf{o}})$, assigning vertices in the same $\omega$-component of $\mathsf{o}-{\mathfrak{e}_{3}}$ the color $\mathsf{blue}$, assigning vertices in the $\omega$-component of $\partial\Lambda_{n,n}^{\prime}$ the color $\mathsf{red}$, and assigning all other $\omega$-components colors chosen independently at random.
@@ Appendix A Generalized cluster expansion
The integral above is a constant, so it and the term $c^{r}(\tfrac{q}{2})^{r/c}$ can be absorbed into the term $e^{-\beta r/c_{2}}$ by increasing the constant $c_{2}$ to some sufficiently large (but $\beta$ independent) constant $K$. ∎
@@
Appendix B Proof of Theorem 2.5
Recall the definitions of walls and ceilings, and their excess areas from Section 3.3. Recall from [14] the notion of groups of walls (collections of nearby walls clustered by a threshold on their interactions through $\mathbf{g}$ relative to their excess areas) that are essential to being able to execute the deletion operations and establish rigidity in $\bar{\pi}^{0}_{\Lambda^{\prime}_{n,\infty}}$.
The proof of Theorem 2.5 with $m=\infty$, found in [14], goes by considering the Peierls-type map that deletes a group of walls $\mathfrak{W}$ from the standard wall representation of $\mathcal{I}_{\mathsf{full}}$, and then bounds the entropy lost from the operation (controlled by enumerating the number of possible groups of walls of size $k$ nesting a vertex $x$ by $C^{k}$) by the energetic gain from the deletion, which is given by $\exp(-\beta\sum_{W\in\mathfrak{W}}\mathfrak{m}(W))$.
When the domain is capped at distance $m$ away from height zero, there is the potential concern of interactions with the hard floor (the extreme case of $m=0$ of course being the same form as the interactions that cause many of the difficulties encountered in this present paper). However, when it is at distance $\Omega(n)$ away in both directions, the interaction is fairly crudely bounded by the energetic cost of reaching up to such a high height.
To begin, notice that for any configuration $\mathcal{I}_{\mathsf{full}}$ having $\overline{\operatorname{ht}}(\mathcal{I}_{\mathsf{full}})=\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\leq m/4$, say, the operation of deleting a group of walls can induce a shift, at most, of $m/4$ on its internal walls, so after the operation, the interface will retain distance at least $m/2$ to the boundary conditions at heights $\pm m$. Then, per Remark 2.4, the only difference in the change of weights caused by this map, relative to the same map in $\Lambda_{n,\infty}^{\prime}$ is an additive $Cn^{2}e^{-n/2}$ from the radius $\mathbf{r}(f,I;f^{\prime},I^{\prime})$ possibly being attained from faces in the boundary. This is of course negligible compared to the cost of $\beta r$ for deletion of a group of wall of excess area $r$.
Thus, without any change to the argument that takes one from Proposition 2.3 to Theorem 2.5, we get for $m\geq 4k$, for every $x$,
$\bar{\pi}^{0}_{\Lambda^{\prime}_{n,m}}(\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\geq k,\,\max_{y}\overline{\operatorname{ht}}_{y}(\mathcal{I}_{\mathsf{full}})\leq m/4)\leq\exp(-a_{p}k)\,$,
It therefore suffices to upper bound the probability of $\max_{y}\overline{\operatorname{ht}}_{y}(\mathcal{I}_{\mathsf{full}})\geq m/4$, for $m\geq n$, by $\exp(-a_{p}^{\prime}m)$ as that will be smaller than $\exp(-a_{p}k)$ up to changing the constant $a_{p}$. For that purpose, consider any interface with $\max_{y}\overline{\operatorname{ht}}_{y}(\mathcal{I}_{\mathsf{full}})\geq m/4$. There must be a nested sequence of groups of walls $\overline{\mathfrak{W}}$ (meaning the union of the groups of walls of a nested sequence of walls $W_{1},...,W_{N}$) whose total excess area satisfies $\mathfrak{m}(\overline{\mathfrak{W}})\geq m/4$.
Consider the map $\Phi$ that iteratively,
  1. (1)
    Picks a nested sequence of groups of walls whose total excess area exceeds $m/4$ (if one exists) and deletes it from the standard wall representation;
  2. (2)
    If no more exist, stops and outputs the resulting interface as $\Phi(\mathcal{I}_{\mathsf{full}})$.
Notice that the resulting interface must be a valid interface in $\Lambda_{n,m}^{\prime}$ with $\overline{\operatorname{ht}}(\Phi(\mathcal{I}_{\mathsf{full}}))\leq m/4$. When considering the weight gain from this map, as compared to the infinite-cylinder analysis, the only difference is the possibility that the radius of congruence $\mathbf{r}(f,\mathcal{I}_{\mathsf{full}};f^{\prime},\Phi(\mathcal{I}_{\mathsf{full}}))$ is attained by interactions between a face $f$ interior of a wall that was deleted, and the boundary conditions at heights $\pm m$. Indeed if $\mathcal{I}_{\mathsf{full}}$ is any such interface, and $\mathfrak{W}$ is the set of walls deleted, then by [14]
$\Bigg{|}\log\frac{\bar{\pi}_{\Lambda^{\prime}_{n,\infty}}^{0}(\mathcal{I}_{\mathsf{full}})}{\bar{\pi}_{\Lambda^{\prime}_{n,\infty}}^{0}(\Phi(\mathcal{I}_{\mathsf{full}}))}+\beta\mathfrak{m}(\mathfrak{W})\Bigg{|}\leq C\mathfrak{m}(\mathfrak{W})\,$, (B.1)
while at the same time, by Proposition 2.3 and Remark 2.4,
$\Bigg{|}\log\frac{\bar{\pi}_{\Lambda_{n,m}^{\prime}}^{0}(\mathcal{I}_{\mathsf{full}})}{\bar{\pi}_{\Lambda_{n,m}^{\prime}}^{0}(\Phi(\mathcal{I}_{\mathsf{full}}))}-\log\frac{\bar{\pi}_{\Lambda^{\prime}_{n,\infty}}^{0}(\mathcal{I}_{\mathsf{full}})}{\bar{\pi}_{\Lambda^{\prime}_{n,\infty}}^{0}(\Phi(\mathcal{I}_{\mathsf{full}}))}\Bigg{|}\leq Kn^{2}e^{-cm/4}+\sum_{f\in\mathsf{Ceil}(\mathfrak{W}):|\operatorname{ht}(f)|\geq\frac{m}{2}}Ke^{-cd(f,\mathcal{L}_{\pm m})}\,.$ (B.2)
where $c,K$ are the constants from Proposition 2.3, and the sum runs over $f\in\mathcal{I}_{\mathsf{full}}$ in ceilings interior to walls of $\mathfrak{W}$. At this point, we wish to bound the number of summands in (B.2) by $\mathfrak{m}(\mathfrak{W})$. Notice that if we partition $A=\rho(\{f\in\mathsf{Ceil}(W):|\operatorname{ht}(f)|\geq m/2\})$ where $\rho(\cdot)$ is used to denote projecting it down to faces of $\mathcal{L}_{0}$, for any $M$ and any partition of $A$ into $\mathfrak{C}_{1},...,\mathfrak{C}_{M}$, we have
$|A|\leq\sum_{i=1}^{M}\operatorname{diam}(\mathfrak{C}_{i})^{2}\leq\sum_{i=1}^{M}\operatorname{diam}(\mathfrak{C}_{i})\cdot n\,$, (B.3)
(because at worst, we are overcounting faces with this bound). On the other hand, for every height between j=1,,m/2j=1,...,m/2, there must have been a collection of wall faces in $\mathfrak{W}$ nesting these ceilings, and therefore at each height, there was some partition $\mathfrak{C}_{1}^{j},...,\mathfrak{C}_{M_{j}}^{j}$ such that those deleted wall faces at height $j$ contributed at least $\sum_{i=1}^{M_{j}}\operatorname{diam}(\mathfrak{C}_{M_{j}}^{j})$. By (B.3), for each $j$, this is at least $|A|/n$ and we sum over j=1,,m/2j=1,...,m/2 to get that $\mathfrak{m}(\mathfrak{W})\geq|A|\frac{m}{2n}$. Since $m\geq n/2$, this is at least $|A|/4$, so the term in (B.2) the right hand side is bounded by $Kn^{2}e^{-cn/8}+K\mathfrak{m}(\mathfrak{W})$, which combined with (B.1) implies the same bound as in that equation for $\bar{\pi}_{\Lambda_{n,m}^{\prime}}^{0}$ up to a change of $C$ to $C^{\prime}$ for another $\beta$ independent constant.
The only other thing left to verify is that the entropy lost by deletion of these walls is controlled by $C^{\mathfrak{m}(\mathfrak{W})}$ for a universal constant $C$. To see this, consider the number of pre-images of a $\mathcal{J}=\Phi(\mathcal{I}_{\mathsf{full}})$, having $\mathfrak{m}(\mathcal{I};\mathcal{J})=L$. We first pick a number $I\in\{1,...,L/m\}$ of iterations of step (1) of the map $\Phi$ before it terminated; then pick $I$ many vertices $v_{1},...,v_{I}$ in $\frac{n}{2},\frac{n}{2}]^{2}\times\{0\}$, allocate the total excess area $L$ between them with each being allocated at least $m/4$ into $L_{1},...,L_{I}$, and then enumerate over nested sequences of groups of walls containing a vertex $v_{i}$ by $C^{L_{i}}$ per [14, Proof of Lemma 15]. In total, this bounds the entropy loss by
$\frac{L}{m}\max_{I\leq L/m}n^{2I}2^{L}\prod_{i=1}^{I}C^{L_{i}}\leq C_{1}^{L}$
where we used the fact that $m\geq n/2$. Since we established that the weight change from such maps with $\mathfrak{m}(\mathcal{I};\mathcal{J})=L$ is $\exp(-(\beta-C)L)$, summing over $L\geq m/4$ is at most $\exp(-(\beta-C^{\prime})n/4)$ concluding the proof. ∎
@@ Acknowledgments
The authors thank the anonymous referees for their careful reading and helpful comments. The research of R.G. is supported in part by NSF DMS-2246780. The research of E.L. was supported by the NSF grant DMS-2054833.